OpenStudy (anonymous):

Andrew invested $20,000 in mutual funds and received a sum of $35,000 at the end of the investment period. Calculate the ROI.

OpenStudy (anonymous):

@mcghee896 assuming ROI = Rate of Interest Principle = $20000 AMount= $350000 Interest = $35000 - $20000 = $15000 Time = PLease specify the time Formula if : \[SI = \frac{ P \times R \times T }{ 100 }\] \[R = \frac{ SI }{ P \times T }\] Note SI = SImple Interest or Interest Forumula if: \[A = P(1 + rt)\] Then \[R = \left(\frac{ 1 }{ T }\right)\left( \frac{ A }{P } - 1\right)\] Both the Formula's are correct...only you need to switch -in the value of Time period T

OpenStudy (anonymous):

thanks man